# Systematic bias from waveform modeling for binary black hole populations in next-generation gravitational wave detectors

### v1

@@ I Introduction
Some of the existing literature estimates the impact of waveform errors on a handful of loud binaries. Owen et al. (2023) demonstrate the possibility of significant systematic biases due to waveform errors, even as early as O4. Pürrer and Haster (2020) compute the expected parameter biases for a few high-SNR binaries in CE/ET to be in excess of $100\sigma$. They show, albeit briefly, that waveform errors could also bias population model parameters. They further estimate that the mismatch between two given waveform models must be improved by $\sim 3$ orders of magnitude to ensure accurate parameter estimation for the loudest binaries in XG detectors. Hu and Veitch (2022) also conclude that a $3-4$ order of magnitude improvement in waveform accuracy would be required at the SNRs expected in future ground-based detectors. Previous work has also shown that systematic bias will play a crucial role for the high-SNR massive black hole mergers observed by the Laser Interferometer Space Antenna (LISA) Berti (2006); Cutler and Vallisneri (2007); Pitte et al. (2023), and that parameterized tests of general relativity using high-SNR detections can produce biased results due to waveform modeling errors (Moore et al., 2021; Saini et al., 2022, 2023; Hu and Veitch, 2023).
@@ III Astrophysical Population of Binary Black Holes
The range of possible redshifts is chosen to be $z\in[0.02,50]$ to match the population in Ref. (Borhanian and Sathyaprakash, 2022). Wherever necessary, the redshift is converted to $D_{\rm L}$ (and vice-versa) using the Planck18 (Aghanim et al., 2020) cosmology. For simplicity, we assume that the star formation redshift distribution is the same as the merger redshift distribution, i.e., we do not explicitly model the delay times. The effect of this assumption will likely be to underestimate the significance of waveform biases, since a more careful modeling of delay times will cause BBHs to merge at lower redshifts, and thus produce smaller statistical errors on average.
@@ IV Results
In our estimates of statistical and systematic parameter estimation errors, we will limit our attention to the loudest BBH sources, i.e., those with SNR $\geq 100$. These systems (roughly 3000 binaries in our population model) are particularly valuable for astrophysical applications due to the small parameter uncertainties, and they are also most affected by any systematic biases between different waveform models. The event posteriors of high-SNR BBHs are also generally well approximated by Gaussian distributions, and so the Fisher information formalism used in this work is expected to be a good approximation for these events. For the same reason, we also expect the systematic bias computed under the linear signal approximation to be reliable for these binaries. Wherever necessary, we impose a limit in post-processing to ensure that the biases do not exceed physical bounds.
@@ V Comparison with full parameter estimation
We find that the biased estimates from Cutler-Vallisneri, $\theta_{i,\rm biased}$, for all parameters relevant to this study are close to the best-estimate parameters recovered with BILBY, whether we do or do not impose priors. The Fisher errors $\sigma_{i,\rm stat}$, instead, tend to be larger than the corresponding intervals in the $1$D marginalized posteriors. For $\mathcal{M}_{\rm c,det}$, $\eta$, and $D_{\rm L}$ (or equivalently $m_{1,\rm det}$, $m_{2,\rm det}$, and $z$) the discrepancy between Fisher errors and $1\sigma$ intervals in the marginalized posteriors is within a factor of $\lesssim 2$. If we do not impose a prior on the the spins, the Fisher errors on $\chi_{1,\rm z}$ and $\chi_{2,\rm z}$ (dashed black contours) are larger by more than an order of magnitude than the corresponding confidence intervals in the $1$D marginalized posteriors from BILBY. When we impose a Gaussian prior (solid black contours), the disagreement decreases to a factor of $\lesssim 3$ for both spin components. However, due to parameter degeneracies, imposing a prior on the spins also affects $\mathcal{M}_{\rm c,det}$ and $\eta$. For $\mathcal{M}_{\rm c,det}$, the Fisher error goes from being $\sim 50\%$ larger than the $1\sigma$ interval in the marginalized posteriors to being $\sim 15\%$ smaller. For $\eta$, the Fisher error goes from $\sim 2$ times larger to $\sim 40\%$ smaller.
@@ Acknowledgements.
The authors are thankful to A. Dhani, S. Völkel, and A. Buonanno for pointing out the importance of waveform alignment, as well as for generously helping us in implementing the alignment procedure in our code. We also thank Ssohrab Borhanian, Ken Ng, Andrea Antonelli, Carl-Johan Haster, Nico Yunes and Bore Gao for helpful discussions. V.K., R.C., L.R. and E.B. are supported by NSF Grants No. AST-2006538, PHY-2207502, PHY-090003 and PHY-20043, by NASA Grants No. 20-LPS20-0011 and 21-ATP21-0010, by the John Templeton Foundation Grant 62840, by the Simons Foundation, and by the Italian Ministry of Foreign Affairs and International Cooperation Grant No. PGR01167. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (rockfish.jhu.edu), which is supported by the NSF Grant No. OAC-1920103.

### v2

@@ I Introduction
Some of the existing literature estimates the impact of waveform errors on a handful of loud binaries. Studies by Owen et al. (2023) and Read (2023), among others, demonstrate the possibility of significant systematic biases due to waveform errors, even as early as O4. Pürrer and Haster (2020) compute the expected parameter biases for a few high-SNR binaries in CE/ET to be in excess of $100\sigma$. They show, albeit briefly, that waveform errors could also bias population model parameters. They further estimate that the mismatch between two given waveform models must be improved by $\sim 3$ orders of magnitude to ensure accurate parameter estimation for the loudest binaries in XG detectors. Hu and Veitch (2022) also conclude that a $3-4$ order of magnitude improvement in waveform accuracy would be required at the SNRs expected in future ground-based detectors. Previous work has also shown that systematic bias will play a crucial role for the high-SNR massive black hole mergers observed by the Laser Interferometer Space Antenna (LISA) Berti (2006); Cutler and Vallisneri (2007); Pitte et al. (2023), and that parametrized tests of general relativity using high-SNR detections can produce biased results due to waveform modeling errors (Moore et al., 2021; Saini et al., 2022, 2023; Hu and Veitch, 2023; Dutta Roy and Saini, 2024).
@@ III Astrophysical Population of Binary Black Holes
The range of possible redshifts is chosen to be $z\in[0.02,50]$ to match the population in Ref. (Borhanian and Sathyaprakash, 2022). Wherever necessary, the redshift is converted to $D_{\rm L}$ (and vice versa) using the Planck18 (Aghanim et al., 2020) cosmology. For simplicity, we assume that the star formation redshift distribution is the same as the merger redshift distribution, i.e., we do not explicitly model the delay times. The effect of this assumption will likely be to underestimate the significance of waveform biases, since a more careful modeling of delay times will cause BBHs to merge at lower redshifts, and thus produce smaller statistical errors on average.
@@ IV Results
In our estimates of statistical and systematic parameter estimation errors, we will limit our attention to the loudest BBH sources, i.e., those with SNR $\geq 100$. These systems (roughly 3000 binaries in our population model) are particularly valuable for astrophysical applications due to the small parameter uncertainties, and they are also most affected by any systematic biases between different waveform models. The event posteriors of high-SNR BBHs are also generally well approximated by Gaussian distributions, and so the Fisher information formalism used in this work is expected to be a good approximation for these events. For the same reason, we also expect the systematic bias computed under the linear signal approximation to be reliable for these binaries. Wherever necessary, we impose a limit in postprocessing to ensure that the biases do not exceed physical bounds.
@@ V Comparison with full parameter estimation
We find that the biased estimates from Cutler-Vallisneri, $\theta_{i,\rm biased}$, for all parameters relevant to this study are close to the best-estimate parameters recovered with BILBY, whether we do or do not impose priors. The Fisher errors $\sigma_{i,\rm stat}$, instead, tend to be larger than the corresponding intervals in the $1$D marginalized posteriors. For $\mathcal{M}_{\rm c,det}$, $\eta$, and $D_{\rm L}$ (or equivalently $m_{1,\rm det}$, $m_{2,\rm det}$, and $z$) the discrepancy between Fisher errors and $1\sigma$ intervals in the marginalized posteriors is within a factor of $\lesssim 2$. If we do not impose a prior on the spins, the Fisher errors on $\chi_{1,\rm z}$ and $\chi_{2,\rm z}$ (dashed black contours) are larger by more than an order of magnitude than the corresponding confidence intervals in the $1$D marginalized posteriors from BILBY. When we impose a Gaussian prior (solid black contours), the disagreement decreases to a factor of $\lesssim 3$ for both spin components. However, due to parameter degeneracies, imposing a prior on the spins also affects $\mathcal{M}_{\rm c,det}$ and $\eta$. For $\mathcal{M}_{\rm c,det}$, the Fisher error goes from being $\sim 50\%$ larger than the $1\sigma$ interval in the marginalized posteriors to being $\sim 15\%$ smaller. For $\eta$, the Fisher error goes from $\sim 2$ times larger to $\sim 40\%$ smaller.
@@ Acknowledgements.
The authors are thankful to A. Dhani, S. Völkel, and A. Buonanno for pointing out the importance of waveform alignment, as well as for generously helping us in implementing the alignment procedure in our code. We also thank Ssohrab Borhanian, Ken Ng, Andrea Antonelli, Carl-Johan Haster, Nico Yunes, and Bore Gao for helpful discussions. V.K., R.C., L.R. and E.B. are supported by NSF Grants No. AST-2006538, No. PHY-2207502, No. PHY-090003, and No. PHY-20043, by NASA Grants No. 20-LPS20-0011 and No. 21-ATP21-0010, by the John Templeton Foundation Grant 62840, by the Simons Foundation, and by the Italian Ministry of Foreign Affairs and International Cooperation Grant No. PGR01167. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (rockfish.jhu.edu), which is supported by the NSF Grant No. OAC-1920103.